OpenStudy (blackops2luvr):

what is 10^-4 written in standard form?

OpenStudy (blackops2luvr):

here are the ancwers: A. 0.00001 B. 0.0001 C. 1,000 D. 10,000

OpenStudy (blackops2luvr):

i think b ic the ancwer :)

OpenStudy (anonymous):

A, I think.

OpenStudy (anonymous):

10^-4 = 1/(10^4) = 1/10000 = .0001